Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

[PR] Problem on Probability

Expert replies
by Xia_devil » Tue Jun 24, 2008 2:04 am
Hey guys

How do i solve this problem ?

A man chooses an outfit from 3 different shirts, 2 different pairs of shoes, and 3 different pants. If he randomly selects 1 shirt, 1 pair of shoes, and 1 pair of pants each morning for 3 days, what is the probability that he wears the same pair of shoes each day, but that no other piece of clothing is repeated?
Last edited by Xia_devil on Tue Jun 24, 2008 2:50 am, edited 1 time in total.
With regards,
Join the discussion
Source: — Problem Solving |

by atlantic » Tue Jun 24, 2008 2:49 am
Hi xia_devil,

Answer is 1/42. Post calculation if correct.
Join the discussion

Re: [PR] Problem on Probability

by Motherjane » Tue Jun 24, 2008 3:45 am
Xia_devil wrote:Hey guys

How do i solve this problem ?

A man chooses an outfit from 3 different shirts, 2 different pairs of shoes, and 3 different pants. If he randomly selects 1 shirt, 1 pair of shoes, and 1 pair of pants each morning for 3 days, what is the probability that he wears the same pair of shoes each day, but that no other piece of clothing is repeated?
The answer is 1/81.

Here's how I went abt it.

First day selection - (3C1/3) * (2C1/2) * (3C1/3) = 1
Second day selection - (2C1/3) * (1/2) * (2C1/3) = 2/9
Third day selection - (1/3) * (1/2) * (1/3) = 1/18

Probability - 1*2/9*1/18 = 1/81 = (1/3) ^ 4
Join the discussion

by atlantic » Tue Jun 24, 2008 5:26 am
Motherjane,

You're right. Again a mistake in my calculation.

Day 1: 1 (because you can pick anything)
Day 2: 1/2 (for the shoes) * 2/3 (for the shirts) * 2/3 (for the pants)
Day 3: 1/2 (for the shoes) * 1/3 (for the shirts) * 1/3 (for the pants)

Then 1*1/2*2/3*2/3*1/2*1/3*1/3= 4/324=1/81.
Join the discussion